Monday 9 August 1999 7.00am EST
|The following is an overview of the events that occurred over the last 24 hours.
ACIS door opened nominally on 8/8/99 at approximately 7:00 PM EDT.
Successfully completed IPS deactivation.
The optical cavity pressure was 8.1E-06 Torr. Pressure remains the same during the last 24 hours. Because of the extended time required for pressure decay, no more attempts to turn on HRC high voltage are planned until after first light.
Events scheduled for the next 24 hours
Return to Status Reports